# Attentive Eraser: Unleashing Diffusion Model’s Object Removal Potential via Self-Attention Redirection Guidance

### v7

@@ Introduction
Figure 2: The overview of our proposed Attentive Eraser which consists of two parts: (a) Attention Activation and Suppression (AAS), a self-attention mechanism modification operation tailored to address the challenges inherent to the object removal task, aims to make the foreground object area’s generation more attentive to the background while erasing the object’s appearance information. Additionally, Similarity Suppression (SS) serves to suppress the heightened attention to similar objects that may arise due to the inherent nature of self-attention. (b) Self-Attention Redirection Guidance (SARG), a guidance method applied in the diffusion reverse sampling process, which utilizes redirected self-attention through AAS to guide the sampling process towards the direction of object removal.
@@
Combining the above analysis, we propose an approach that is both simple and effective: AAS (as shown in Figure 2(a)). Activation refers to increasing $A_{l,t}^{obj\rightarrow bg}$, which serves to enhance the attention of the foreground-generating region to the background. In contrast, Suppression refers to decreasing $A_{l,t}^{obj\rightarrow obj}$ and $A_{l,t}^{bg\rightarrow obj}$, which entails the suppression of the foreground region’s information about its appearance and its effect on the background. Given the intrinsic characteristics of the Softmax function, AAS can be simply achieved by assigning $S_{l,t}^{obj\rightarrow obj}$ to $\infty$- ∞, thereby the original semantic information of the foreground objects is progressively obliterated throughout the denoising process. In practice, the aforementioned operation is achieved by the following equation:
@@
[…]
@@ A. Detailed analysis of SARG
where the pre-trained score estimation network $\epsilon_{\theta}$ and the AAS processed network $\text{AAS}(\epsilon_{\theta})$ are approximations of $-\sigma_{t}\nabla_{x_{t}}\log p\left(x_{t}\mid y\right)$ and $-\sigma_{t}\nabla_{x_{t}}\log p\left(x_{t}\mid y_{ASS}\right)$, respectively.
@@ C. Additional Experimental Details
1:  $x_{0}\leftarrow\text{VAE-Encoder}\left(X_{0}\right)$.
2:  $x_{T_{I}}\leftarrow\sqrt{\bar{\alpha}_{T_{I}}}x_{0}+\sqrt{\left(1-\bar{\alpha}_{T_{I}}\right)}\epsilon$.
3:  $z_{T_{I}}\leftarrow x_{T_{I}}$.
@@
During the evaluation process, aside from Inst-Inpaint requiring 256$\times$256 image inputs, we utilize 512$\times$512 images as the input for each method. When calculating the metrics, all output images are resized to 512$\times$512. Given the necessity for the CLIP consensus metric to assess image generation across various random seeds and subsequently calculate the standard deviation of the CLIP embeddings within the foreground object region. We extract 6000 images from the testing dataset of 10000 and generate corresponding results using the random seeds (123, 321, 777) for the computation of this metric. The remaining metrics are calculated using results from the testing dataset of 10000 images generated with the random seed 123.

### v8

@@ Introduction
Figure 2: The overview of our proposed Attentive Eraser which consists of two parts: (a) Attention Activation and Suppression (AAS), a self-attention mechanism modification operation tailored to address the challenges inherent to the object removal task, aims to make the foreground object area’s generation more attentive to the background while erasing the object’s appearance information. Additionally, Similarity Suppression ($S$) serves to suppress the heightened attention to similar objects that may arise due to the inherent nature of self-attention. (b) Self-Attention Redirection Guidance (SARG), a guidance method applied in the diffusion reverse sampling process, which utilizes redirected self-attention through AAS to guide the sampling process towards the direction of object removal.
@@
Combining the above analysis, we propose an approach that is both simple and effective: AAS (as shown in Figure 2(a)). Activation refers to increasing $A_{l,t}^{obj\rightarrow bg}$, which serves to enhance the attention of the foreground-generating region to the background. In contrast, Suppression refers to decreasing $A_{l,t}^{obj\rightarrow obj}$ and $A_{l,t}^{bg\rightarrow obj}$, which entails the suppression of the foreground region’s information about its appearance and its effect on the background. Given the intrinsic characteristics of the Softmax function, AAS can be simply achieved by assigning $S_{l,t}^{obj\rightarrow obj}$ to -$\infty$, thereby the original semantic information of the foreground objects is progressively obliterated throughout the denoising process. In practice, the aforementioned operation is achieved by the following equation:
@@
SD2.1inp
3.805 0.3012 8.852 0.1143 21.89
SD2.1inp
4.019 0.3083 7.194 0.1209 22.27
PowerPaint
6.027 0.2887 10.02 0.0984 22.74
Inst-Inpaint
11.42 0.4095 43.47 0.0913 23.02
LAMA
7.533 0.2189 6.091 - 23.57
SD2.1+SIP w/o SARG
5.98 0.2998 15.58 0.1347 22.05
SD2.1+SIP w/ SARG(ours)
7.352 0.3113 5.835 0.0734 23.56
SD2.1+DIP w/ SARG(ours)
7.012 0.2995 5.699 - 23.43
@@ A. Detailed analysis of SARG
where the pre-trained score estimation network $\epsilon_{\theta}$ and the AAS processed network $\text{AAS}(\epsilon_{\theta})$ are approximations of $\sigma_{t}\nabla_{x_{t}}\log p\left(x_{t}\mid y\right)$ and $\sigma_{t}\nabla_{x_{t}}\log p\left(x_{t}\mid y_{ASS}\right)$, respectively.
@@ C. Additional Experimental Details
[…]
@@
During the evaluation process, aside from Inst-Inpaint requiring $\times256$ image inputs, we utilize $\times512$ images as the input for each method. When calculating the metrics, all output images are resized to $\times512$. Given the necessity for the CLIP consensus metric to assess image generation across various random seeds and subsequently calculate the standard deviation of the CLIP embeddings within the foreground object region. We extract 6000 images from the testing dataset of 10000 and generate corresponding results using the random seeds (123, 321, 777) for the computation of this metric. The remaining metrics are calculated using results from the testing dataset of 10000 images generated with the random seed 123.